Cesars Bottlers bottles soft drinks in a factory that can operate either one shift, two shifts, or three shifts per day. Each shift is eight hours long. The factory is closed on weekends. The sales price of $4 per case bottled and the variable cost of $2.7 per case remain constant regardless of volume Cesars Bottlers can increase volume by opening and staffing additional shifts. The company has the following three choices Daily Volume Range Total Fixed Costs per (Number of Cases Bottled) Day 1 Shift 2 Shifts 3 Shifts (0-2,000) (2,001-3,600) (3,601-5,000) $ 2,400 4,480 6,140 Required a. Calculate the break-even point(s). (Round up your answers to the nearest whole number.) Break-even points 1 shift 2 shifts 3 shifts cases cases cases b-1. Calculate the profit (or loss) for each alternative, assuming Cesar can sell all the cases he can bottle. (Loss amounts should be indicated witha minus sign.) Profit (Loss) 1 shift 2 shifts 3 shifts b-2. Should Cesar operate one, two, or three shifts? O 1 shift O 2 shifts O 3 shifts
Answer #1

BEP- FIXED EXPENSES/UNIT CONTRIBUTION. ONE SHIFT- 2400/4-2.7= 1846 . two shifts-4480/4-2.7= 3446. Three shifts - 6140/4-2.7= 4723 units. 2. Profit- one shift - sales- cost= 2000*(4-2.7)-2400=200, two shifts= 3600*(4-2.7)-4480=200. Three shifts= 5000*(4-2.7)-6140= 360. I will select three shifts because it generate more profits.